OpenStudy (lucaz):

find the values ​​of x for which the expression is positive: x^2-1/x^2-3x

OpenStudy (fifciol):

\[\frac{ x^2-1 }{ x^2-3x }>0 \]when \[(x^2-1)(x^2-3x)>0\]

OpenStudy (lucaz):

how did you find that?

OpenStudy (lucaz):

lol, I see

OpenStudy (fifciol):

so x(x-1)(x+1)(x-3)>0  so \[x \in (-\infty; -1) \cup (0;1) \cup (3;\infty)\] it's the property of rational numbers